OpenStudy (abbycross167):

can someone please help me answer this question? I could really use some help... Question~ Solve the systems using the Substitution method. -3x + y = 0 3x - y = -12

OpenStudy (mathstudent55):

To use the substitution method, the first thing you need to do is to solve one equation for one variable.

OpenStudy (mathstudent55):

Use the first equation. Can you solve it for y?

OpenStudy (mathstudent55):

Since you have -3x on the left side, but you want y alone, add 3x to both sides. What do you get?

OpenStudy (abbycross167):

-3x + y = 0-3x? y=0?

OpenStudy (mathstudent55):

No. You need to add 3x to both sides. -3x + y = 0 Add 3x to both sides: -3x + 3x + y = 0 + 3x y = 3x

OpenStudy (mathstudent55):

Now you have the first equation solved for y. Now we do the substitution step. Since y = 3x, where we see y in the second equation, we substitute y with 3x.

OpenStudy (abbycross167):

oo ok I'm sorry I thought I had to subtract

OpenStudy (mathstudent55):

Here is the second equation. 3x - y = -12 We see y on the left side. We now rewrite the equation, substituting y with 3x. 3x - 3x = -12 You see that y was replaced by 3x?

OpenStudy (abbycross167):

yes sir/ma'am

OpenStudy (mathstudent55):

Now we simplify the left side. Since both terms on the left side have only the variable x, they are like terms. We can combine them together. 3x - 3x = 0, so we write: 3x - 3x = -12 0 = -12 Now we got to a false statement of equality. 0 in not equal to -12. Since after doing several steps correctly we got to a false statement, that means the system of equations has no solution.

OpenStudy (mathstudent55):

Answer: no solution

OpenStudy (abbycross167):

oh ok, thank you soo much!! you're a great help!

OpenStudy (mathstudent55):

You're welcome.